OpenStudy (anonymous):

how do you make this equation in slope-intercept form. x/2- y + 6 = 0

OpenStudy (jackellyn):

slope intercept form y=mx +b x/2 -y +6 =0 Subtract 6 from both sides x/2 -y = -6 Multiply 2 on both sides x -y = -6(2) Subtract x from both sides -y = -12 -x Divide by -1 on both sides y = 12 + x y = x + 12

OpenStudy (anonymous):

which one would it be out of these? A. y = - x/2- 6 B.y =x/2 + 6 C.y =x/2 - 6

OpenStudy (jackellyn):

-y = -6 -x/2 y= 6 + x/2 y = x/2 +6 They subtracted the entire (x/2) on both sides

OpenStudy (anonymous):

ohhh okkk thnxx

OpenStudy (ajprincess):

When u multiply by 2 both sides it becomes x-2y=-12 Nt x-y=-12 @jackellyn.

OpenStudy (jackellyn):

you're right @ajprincess

OpenStudy (anonymous):

iunderstand now i appreciate your consideration

OpenStudy (jackellyn):

x-2y=-12 -2y = -12 -x y = 6 + x/2 If i had done that step right I would've gotten the same answer.

OpenStudy (ajprincess):

yes u r right.